OpenStudy (anonymous):

Britain bought tickets at a local carnival. He rode the carousel 5 times and the ferris wheel 2 times. Each ride required 3 tickets. Which of the following represents the total number of tickets Britain used? 3 times the quantity of 5 plus 2 5 times the quantity of 2 plus 3 3 times 5 times 2 5 plus 2 times 3

OpenStudy (anonymous):

i would go with the first one

OpenStudy (anonymous):

total number of rides is \(5+2\) and each was 3 tickets so \(3(5+2)\) what a weird question!!

OpenStudy (anonymous):

At this rate i wont graduate

OpenStudy (anonymous):

5 rides and 2 rides is 7 rides, 3 tickets for each so total number of tickets is \(7\times 3=21\)

OpenStudy (anonymous):

must be some new math i never seen

OpenStudy (anonymous):

yeah if they keep asking you questions like these, with the proposed answers, you could do this forever before you came up with the 21 tickets you could have computed instantly in your head like i said, new math...

OpenStudy (anonymous):

wait, let me guess home school packet? (not to be critical)

OpenStudy (anonymous):

I take my college classes online

OpenStudy (anonymous):

i see

OpenStudy (agent0smith):

It is a bit odd, but i guess it's testing the ability to interpret word problems and have answers in word form... and maybe testing the distributive property. You could do it this way each ride reqs 3 tickets: rode the carousel 5 times = 5*3 ferris wheel 2 times = 2*3 total number of tickets = 5*3 + 2*3 = 3(5+2)